# Eye-tracked Virtual Reality: A Comprehensive Survey on Methods and Privacy Challenges

### v2

@@ IV-A Hardware, Datasets, and Algorithms
In this section, we first provide insights into the works about hardware and datasets that are relevant to eye-tracked VR in Sections IV-A1 and IV-A2, respectively. Then, in Sections IV-A3IV-A4, and IV-A5, we proceed to discuss eye-tracking-related algorithms for eye region segmentation, gaze estimation, and eye event classification, respectively. We briefly summarize the papers discussed in this section in Table I. Eye tracking in immersive VR using HMDs is essentially different from other real-world scenarios, as users mostly wear HMDs on their heads, and sensors only see the eye regions rather than the whole face. Therefore, if not explicitly mentioned, we reduce the focus to such setups in this work. For an extended evaluation of eye tracking, such as gaze estimation models regardless of immersive VR or real-world scenarios, we refer the reader to the survey papers by Cheng et al. [27] for appearance-based gaze estimation with deep learning, by Akinyelu and Blignaut [28] for convolutional neural network (CNN)-based gaze estimation, and by Kar and Corcoran [29] for gaze estimation in consumer platforms.
TABLE I: A non-exhaustive overview of papers discussed in Section IV-A.
@@
[32, 33] SMI add-on eye tracker + HTC Vive / Oculus Rift DK2
[34, 35] HTC Vive Pro Eye, Varjo VR-1, Fove-0
@@
[59] Using CycleGANs for segmentation, refinement, and generation
@@
[62] Eye-MMS: Using multi-scale inter-connected CNN, lightweight
[63] EyeSeg: Using generalized dice loss function, lightweight
[64] RITnet: Using U-Net + DenseNet, lightweight
@@ IV-A1 Hardware
We begin by examining the hardware necessary for implementing eye-tracked VR. Formerly, most VR headsets (e.g., Oculus Rift [79]) did not deliver built-in eye-tracking functionality. Practitioners needed add-on eye-tracking extensions (e.g., Pupil Labs [80], SensoMotoric Instruments (SMI), which was acquired by Apple [81]) to approach eye tracking on these devices. As eye tracking and VR both grow at an astounding rate, an increasing number of HMDs like HTC Vive Pro Eye [82], Fove-0 [83], and Varjo XR-3 [84] support eye tracking on their own. An example of an HMD device, the Varjo XR-3, which comes with an integrated eye tracker, is shown in Figure 4, with eye images and a part of a sample raw data. Meanwhile, a growing body of research highlights the importance of evaluating the performance of built-in and add-on eye trackers for HMD devices. For instance, Lohr et al. [32] measured the quality of eye data acquired with the SMI add-on eye tracker in the HTC Vive from multiple facets, including spatial accuracy, spatial precision, temporal precision, linearity, and crosstalk. The quality of eye-tracking data recorded by the SMI extension has also been analyzed by Roth et al. [33] in combination with Oculus Rift DK2 [85] regarding tracking precision and fixation accuracy in the context of foveated rendering.
Adhanom et al. [30] reported an average eye-tracking accuracy of 1.23 and a root mean square (RMS) precision of 0.62 on the HTC Vive Pro Eye with their newly developed Unity package called GazeMetrics. In a recent work by Schuetz and Fiehler [31], an average accuracy of 1.08 and mean standard deviation (SD)/RMS precisions of 0.36/0.2 were reported on the same device after outlier correction. The authors also found a significant decrease in accuracy and precision when participants wore vision correction glasses, while the effect of contact lenses was more elusive. Stein et al. [35] conducted a comparison between Fove-0, Varjo VR-1 [86], and HTC Vive Pro Eye using another two metrics: eye-tracking delay and latency. With a delay of 15-52 ms and a latency of 45-81 ms, the Fove-0 outperformed the other two HMDs to a large extent. The assessment of eye-tracked HMD can also go beyond statistical performance. For instance, Maraj et al. [34] compared the HTC Vive Pro Eye with the Varjo VR-1 in three eye-tracking scenarios. The authors reported the comparison results in non-parametric forms that covered immersion, simulation sickness, and visual discomfort.
@@
While add-on eye-tracking extensions and built-in tracking modules have become more commercially affordable, they can still be expensive due to additional hardware and software requirements. To address this issue, Stengel et al. [36] proposed an implementation utilizing dichroic mirrors and a personalizable lens positioning system at an estimated cost of merely \$450. The authors deployed dichroic mirrors to support eye-tracking cameras outside of the field of view (FOV), whereas the lens locating module accounts for different interocular distances. With a model-based gaze estimation algorithm, the system achieved a gaze angle error of 0.5-3.5.
Greenwald et al. [37] suggested an alternative, more affordable system, which uses a smartphone and a Google Cardboard [87] that costs approximately \$15. In this system, gaze estimation is approached through capturing Purkinje images (i.e., reflections of on-screen images on the eyes) using the smartphone’s front-facing camera, and an average error of 5.0 can be achieved. In parallel to them, Ahuja et al. [38] presented a similar framework named EyeSpyVR, which also uses the smartphone selfie camera and a \$10 VR Box headset to achieve coarse gaze tracking. Having tested the framework with 70 subjects, the authors reported an average gaze estimation error of 10.8 and 12.9, respectively, for the use case with and without calibration. Stratmann et al.  [39] deployed another system on Google Cardboard called EyeMR. In contrast to the systems above that are realized using the smartphone selfie camera, EyeMR uses USB cameras and extended IR-LED circuit boards to function gaze estimation and supports both mono- and binocular eye tracking.
@@
Common camera-based eye-tracking systems fall into the category of video-oculography (VOG). VOG-based methods often struggle to find an optimal balance among factors like power consumption, computational cost, latency, and accuracy. There also exist other types of eye trackers, which can be categorized into photosensor-oculography (PSOG), electro-oculography (EOG), and scleral search coil (SSC) according to Duchowski et al. [88] and Rigas et al. [89]. PSOG is akin to VOG because they both often require light sources and optical sensors to capture reflections. In contrast, PSOG usually uses a sparse grid of photodiodes and photosensors, recording significantly fewer pixels than VOG. As a result, PSOG can outperform VOG in terms of sampling rate, computational cost, and power consumption. The works by Li et al. [40] and by Katrychuk et al. [41] fall into this category. In comparison to VOG and PSOG, EOG-based methods, such as by Shimizu and Chernyshov [42] and by Bernal et al. [43], as well as SSC-based methods like by Whitmire et al. [44], commonly put no requirement on light emitters or receivers. The former approaches eye tracking by placing electrodes around the eyes and then measuring voltage changes during eye movement, whereas the latter realizes gaze estimation by requiring users to wear wired contact lenses and then tracking the lenses in a magnetic field. As stated by Adhanom et al. [18], EOG-based methods can be easily implemented in HMDs and can track eyes even when closed, although less accurately. In contrast, SSC-based methods provide excellent accuracy but are often hard to deploy. Beyond the aforementioned methods, eye tracking may also be approached with special devices and algorithms. One example would be the work by Shenoy et al. [90] where the authors achieved gaze estimation using an adaptive optics scanning laser ophthalmoscope, which is a device that can image the retina at high resolution and high frame rate. The authors modeled eye tracking as a joint estimation of retina motion and appearance similar to simultaneous localization and mapping (SLAM [91]). They achieved an accuracy of below $\frac{1}{60^{\circ}}$ at 1 kHz sampling rate.
@@ IV-A2 Datasets
While different types of eye-tracking hardware are essential to obtain eye movement data, researchers often publish publicly available eye-tracking datasets to facilitate benchmarking of data-driven methods, and we discuss these in the following.
@@
Data quality, sample size, and context-dependent information are very important for the performance and robustness of the gaze estimators. To this end, Kim et al. [47] proposed two datasets satisfying such criteria, including a synthetic one using anatomically informed eye and face models and a real-world dataset from 35 participants for near-eye gaze estimation. The authors showed that their trained CNNs perform with accuracy losses of $2^{\circ}$ and $0.5^{\circ}$ for person-independent and personalized setups, respectively. Garbin et al. [45, 46] introduced a well-known dataset called OpenEDS that can be used for not only gaze estimation but also eye region segmentation. The dataset was collected from 152 participants in an immersive virtual environment and covers diverse data types, including pixel-level annotations of eye regions, unlabeled eye images, video sequences, and point clouds.
Figure 4: Left: Varjo XR-3, a high-end HMD device for both VR and AR. Upper right: eye images taken by cameras inside Varjo XR-3. Lower right: some eye features recorded by Varjo XR-3.
@@
In most cases, visual attention is raised in a bottom-up pattern driven by visual stimuli during data collection. Jiang et al. [48] presented a novel dataset called IQVA (i.e., immersive question-directed visual attention) in which tasks drove visual attention in a top-down style. The dataset consists of eye-tracking data from 975 HMD video clips with questions, each annotated by 14 participants. The authors also incorporated a neural network to predict the correctness of attention maps. Agtzidis et al. [49] published an eye-tracking dataset for 360 videos collected from 13 observers together with a novel two-stage eye event annotation pipeline and a rule-based eye event classifier for automatic event labeling. Fuhl et al. [50] introduced the largest ever unified public eye dataset, gathered using seven different head-mounted devices in the real world, VR (54 participants), and AR. The dataset covers various types of eye data, such as 2D and 3D landmarks, eye segmentation information, and 3D eyeball annotation.
More recently, Aziz et al. [51] published a new dataset called SynchronEyes, in which eye movements of 15 participants were recorded simultaneously with a stationary and a wearable eye tracker. Further, Jin et al. [52] presented a new quantitative taxonomy for VR videos that relies on three metrics: camera motion, video quality, and dispersion of region of interest (ROI). The authors created a head and gaze behavior dataset based on this taxonomy from 100 participants and 27 videos.
@@
There are also other multi-modal datasets available. An example would be the OpenNEEDS dataset by Emery et al. [53], which includes 44 participants exploring two virtual spaces, and it contains data of not only eyes but also head, hands, and scenes. Tabbaa et al. [54] introduced a dataset called VREED for emotion recognition where data from different modalities, including eye movement, electrocardiogram (ECG), and galvanic skin response (GSR), were recorded from 34 subjects while watching 360 videos. Recently, Zhang et al. [55] published a new multi-modal dataset called EgoBody, including eye gaze, head motion, and gesture data. At the same time, the work emphasizes human body reconstruction from an egocentric view in VR. Chen et al. [56] also presented a multi-modal dataset called HE-Gaze recently, which was captured from 15 participants using an AR HMD. In addition to eye movement data, the dataset also contains head movement data and near-eye images for both eyes. The authors also introduced an appearance-based gaze estimation algorithm called HE-Tracker that utilizes head movement and can approach eye tracking in real-time at 48 Hz with an accuracy of 3.655.
@@
Collecting high-quality eye-gaze data can be costly and time-consuming. This has led to the development of synthetic 3D models such as UT Multi-View [92], SynthesEyes [93], and UnityEyes [94], as well as generative algorithms [95] and data augmentation approach like gaze redirection [96]. While few of these works are specifically tailored for HMD use, some concepts can be adapted for eye data generation in immersive virtual environments.
@@
An essential step during and after data collection is data annotation, which is often costly. Xue et al. [57] addressed the issue of collecting emotion annotations with eye-tracking data in VR environments. Typically, emotional feedback during the VR experience is annotated retrospectively. As a consequence, emotion labeling can be discrete and time-consuming. By leveraging their previously introduced peripheral visualization frameworks HaloLight and DotSize [97], the authors achieved real-time continuous emotion annotation during data collection from 32 participants and provided temporally precise labels for downstream tasks.
@@ IV-A3 Eye Region Segmentation
Having broadly viewed hardware and datasets for eye-tracked VR, we now provide an overview of algorithms related to gaze estimation on HMDs, beginning with eye region segmentation. Semantic segmentation of the eye region is crucial for gaze estimation and is often the starting point for many gaze estimation algorithms. According to Shen et al. [61], early eye region segmentation algorithms like [98, 99] were mainly driven by iris texture and sclera extraction. Nowadays, an increasing shift towards multi-class (multi-region) eye segmentation that is typically realized through deep learning. Unlike other semantic segmentation tasks, eye region segmentation is often challenging due to low-resolution images and possible blurs.
Kansal and Nathan [58] presented an efficient encoder-decoder network called EyeNet that addresses these issues by using residual blocks [100] in both encoder and decoder to improve gradient flow and using convolutional block attention modules (CBAM [101]) to enhance boundary sharpness and accuracy. The network accomplishes a total score of 0.974 on the EDS evaluation metric. Fuhl et al. [59] explored the applicability of CycleGANs [102] for eye segmentation and suggested three different generative adversarial networks (GANs) for segmentation, image refinement, and image generation, respectively. In particular, these models were trained with cyclic loss to prevent the discriminator from overfitting. Conventionally, eye segmentation is vulnerable to occlusion caused by eyelids and eyelashes. Kothari et al. [60] addressed this challenge with a segmentation framework called EllSeg that is robust against occlusions and can be implemented jointly with other pupil and iris ellipse segmentation methods in an encoder-decoder pattern.
@@
Computational resources, such as computational power, memory, and storage, are often bottlenecks of HMDs. Boutros et al. [62] alleviated the computation burden on embedded systems with their multi-scaled segmentation network (Eye-MMS) based on a multi-scale interconnected CNN. The authors achieved a 3% loss in accuracy compared to the original model while successfully reducing the number of model parameters from 6,574k to 80k. Perry and Fernandez [63] proposed a lightweight encoder-decoder segmentation network named EyeSeg, which adopts a customized loss function combining categorical cross entropy and generalized dice loss function (GDL [103]) to tackle the shortage of labeled data. EyeSeg contains only 190k parameters while yielding a 94.5% mean intersection over union (mIOU) on the 2020 OpenEDS dataset. Another lightweight segmentation network, RITnet [64], combines U-Net [104] and DenseNet [105], which is only 0.98 MB and allows eye segmentation at 300 Hz in real-time, with an mIOU of 95.3%, on the OpenEDS dataset.
@@
Another major challenge in eye region segmentation is the lack of high-quality eye images. Damer et al. [65] overcame this challenge by generating eye images from semantic segmentation with their novel model called D-ID-NET. The pipeline of this model consists of two phases: in the first phase, a domain network (D-Net) synthesizes identity-irrelevant images from semantic labels, and in the second phase, identity-specific information is introduced into the images by an identity-specific network (ID-Net). Both networks are CNNs with the same structure but different training strategies. Buehler et al. [66] implemented a similar concept with their model Seg2Eye, which generates content-preserving eye images from semantic segmentation. The work resembles style transfer [106] in the sense that semantic segmentation defines the content of generated images. In contrast, their styles are controlled by style features extracted from images of the target person.
Shen et al. [61] extensively studied domain adaptation for eye segmentation when only a few source images are annotated while most data are not labeled. They systematically investigated the impact of annotated data by training the model in supervised, unsupervised, and semi-supervised manners, respectively, and varying the amount of labeled eye images in the target domain during training. More recently, Lu et al. [67] proposed an image-to-image translation network called GeoMaskGAN that accounts for geometric consistency. The network consumes a pair of eye images and an eye segmentation mask as input, and it outputs a new pair while reducing the geometric gap between translated images and original ones.

### v3

@@ IV-A Hardware, Datasets, and Algorithms
We first provide insights into the works on hardware and datasets related to eye-tracked VR in Sections IV-A1 and IV-A2, respectively. Then, in Sections IV-A3IV-A4, and IV-A5, we discuss algorithms for eye region segmentation, gaze estimation, and eye event classification, respectively. Table I briefly summarizes the papers in this section. Eye tracking in immersive VR using HMDs is fundamentally different from real-world scenarios, as users typically wear HMDs on their heads, and sensors detect the eye regions rather than the entire face. Therefore, unless explicitly mentioned, we focus on such setups. For an extended evaluation of eye tracking, such as gaze estimation, regardless of immersive VR or the real world, we refer the reader to the surveys on appearance-based gaze estimation with deep learning [21], convolutional neural network (CNN)-based gaze estimation [22], and gaze estimation in consumer platforms [23].
TABLE I: A broad overview of papers in Section IV-A.
@@
[25, 26] SMI add-on eye tracker + HTC Vive / Oculus Rift
[27] HTC Vive Pro Eye, Varjo VR-1, Fove-0
@@
[49] Using CycleGANs for segmentation and generation
@@
[52] Eye-MMS: Using multi-scale inter-connected CNN
[53] EyeSeg: Using generalized dice loss function
[54] RITnet: Using U-Net + DenseNet
@@ IV-A1 Hardware
Most VR headsets (e.g., Oculus Rift [69]) formerly did not deliver built-in eye trackers. Practitioners needed add-on eye trackers (e.g., Pupil Labs [70], SensoMotoric Instruments (SMI), acquired by Apple [71]) to enable eye tracking functionality. As eye tracking and VR grow at an astounding rate, an increasing number of HMDs like HTC Vive Pro Eye [1] and Varjo XR-3 [72] support eye tracking on their own. Figure 3 illustrates an example of an HMD, the Varjo XR-3, which features an integrated eye tracker, with eye images and a portion of sample raw data. Meanwhile, a growing body of research highlights the importance of evaluating the performance of built-in and add-on eye trackers for HMDs. For instance, [25] measured data quality acquired with the SMI eye tracker in the HTC Vive from multiple facets, including spatial accuracy and precision, temporal precision, linearity, and crosstalk. Eye-tracking data quality from an SMI extension has also been analyzed [26], considering tracking precision and fixation accuracy in a foveated rendering context.
An average eye-tracking accuracy of 1.23 and a root mean square (RMS) precision of 0.62 were also reported on the HTC Vive Pro Eye with a Unity package called GazeMetrics [15]. In a recent work [24], an average accuracy of 1.08 and mean standard deviation (SD)/RMS precisions of 0.36/0.2 were reported on the same device after outlier correction. The authors also found a significant decrease in accuracy and precision when participants wore vision correction glasses, while the effect of contact lenses was more elusive. [27] conducted a comparison between Fove-0, Varjo VR-1, and HTC Vive Pro Eye using two other metrics: eye-tracking delay and latency. With a delay of 15-52 ms and a latency of 45-81 ms, Fove-0 outperformed the other HMDs.
@@
While add-on or built-in eye trackers have become more affordable, they can still be expensive due to hardware and software requirements. To address this, [28] proposed an implementation utilizing dichroic mirrors and a personalizable lens positioning system at an estimated cost of \$450. The authors deployed dichroic mirrors to support eye-tracking cameras outside of the field of view (FOV), whereas the lens locating module accounts for different interocular distances. With a model-based gaze estimation algorithm, an angular error of 0.5-3.5 was achieved.
Figure 3: Left: Varjo XR-3, a high-end HMD device for XR. Upper right: eye images taken within the HMD. Lower right: some eye features recorded.
An affordable, alternative system that utilizes a smartphone and a Google Cardboard [73], costing approximately \$15, also exists [29]. In this system, gaze estimation is achieved by capturing Purkinje images (i.e., reflections of on-screen images on the eyes) using the smartphone’s front camera, resulting in an average error of 5.0. In parallel, [30] presented a similar framework named EyeSpyVR, which also uses the smartphone selfie camera and a \$10 VR Box headset to achieve coarse gaze tracking. With 70 subjects, the authors reported an average gaze estimation error of 10.8 and 12.9, respectively, with and without calibration. [31] deployed another system on Google Cardboard called EyeMR. In contrast to systems realized using smartphone selfie cameras, EyeMR utilizes USB cameras and extended IR-LED circuit boards to perform gaze estimation, supporting both monocular and binocular eye tracking.
@@
Standard camera-based eye-tracking systems fall into the category of video-oculography (VOG). VOG-based methods often struggle to strike an optimal balance among factors such as power consumption, computational cost, latency, and accuracy. Other types of eye trackers also exist, which can be categorized into photosensor-oculography (PSOG), electro-oculography (EOG), and scleral search coil (SSC) [74, 75]. PSOG is akin to VOG because they both often require light sources and optical sensors to capture reflections. In contrast, PSOG usually uses a sparse grid of photodiodes and photosensors, recording significantly fewer pixels than VOG. As a result, PSOG can outperform VOG in terms of sampling rate, computational cost, and power consumption [32, 33]. In comparison to VOG and PSOG, EOG-based methods [34, 35], as well as SSC-based methods [36], commonly put no requirement on light emitters or receivers. The former approaches eye tracking by placing electrodes around the eyes and then measuring voltage changes during eye movement, whereas the latter realizes gaze estimation by requiring users to wear wired contact lenses and tracking the lenses in a magnetic field. EOG-based methods can be easily implemented in HMDs and can track eyes, even when closed, although with less accuracy [11]. In contrast, SSC-based methods provide excellent accuracy but are often challenging to deploy. Eye tracking may also be approached with specialized devices and algorithms. One example is achieving gaze estimation using an adaptive optics scanning laser ophthalmoscope, a device that can image the retina at high resolution and high frame rate [76]. This approach models eye tracking as a joint estimation of retina motion and appearance, similar to simultaneous localization and mapping (SLAM [77]), achieving an accuracy of below $\frac{1}{60^{\circ}}$ at 1 kHz sampling rate.
@@ IV-A2 Datasets
While different types of eye-tracking hardware are essential for obtaining eye movement data, researchers often publish publicly available eye-tracking datasets to facilitate the benchmarking of data-driven methods. Data quality, sample size, and context-dependent information are essential for the performance and robustness of the gaze estimators. To this end, [39] proposed two datasets satisfying such criteria, including a synthetic one using anatomically informed eye and face models and a real-world dataset from 35 participants for near-eye gaze estimation, and showed that their trained CNNs perform with accuracy losses of $2^{\circ}$ and $0.5^{\circ}$ for person-independent and personalized setups, respectively. [37, 38] introduced the OpenEDS dataset, which was collected from 152 participants in an immersive virtual environment, for both gaze estimation and eye region segmentation. The dataset covers a diverse range of data types, including pixel-level eye region annotations, unlabeled eye images, video sequences, and point clouds.
@@
In most cases, visual attention is raised in a bottom-up pattern driven by visual stimuli during data collection. [40] presented a novel dataset called IQVA (i.e., immersive question-directed visual attention) in which tasks drove visual attention in a top-down style. The dataset consists of eye-tracking data from 975 HMD video clips, each annotated by 14 participants. [41] published an eye-tracking dataset for 360 videos collected from 13 observers together with an eye event annotation pipeline and a rule-based eye event classifier for event labeling. [42] introduced the largest unified public eye dataset, gathered using seven different HMDs in the real world, VR, and AR, from 54 participants. The dataset encompasses various types of eye data, including 2D and 3D landmarks, eye segmentation information, and 3D eyeball annotations. Further, [43] presented a new taxonomy for VR videos that relies on three metrics: camera motion, video quality, and dispersion of region of interest (ROI), along with a dataset on head and gaze behaviors from 100 participants and 27 videos.
@@
Multimodal datasets also exist, such as the OpenNEEDS [44], which includes 44 participants exploring two virtual spaces and contains data from not only the eyes but also the head, hands, and scenes. Furthermore, the VREED dataset for emotion recognition [45] comprises data from various modalities, including eye movement, electrocardiogram (ECG), and galvanic skin response (GSR), which were recorded from 34 subjects while watching 360 videos. Another multimodal dataset, EgoBody, including eye gaze, head motion, and gesture data [46], also emphasizes human body reconstruction from an egocentric view in VR.
@@
Collecting high-quality eye-tracking data can be costly and time-consuming. This has led to the development of synthetic 3D models such as UT Multi-View [78], SynthesEyes [79], and UnityEyes [80], as well as generative algorithms [81] and data augmentation approaches [82]. While few of these works are specifically tailored for HMDs, some concepts can be adapted for eye data generation in immersive VR.
@@
An essential and costly step during and after data collection is data annotation. [47] studied collecting emotion annotations with eye-tracking data in VR, which is typically annotated retrospectively and can be discrete and time-consuming. By leveraging previous visualization frameworks, HaloLight and DotSize [83], the authors achieved real-time continuous emotion annotation during data collection and provided temporal labels for downstream tasks.
@@ IV-A3 Eye Region Segmentation
This section provides an overview of algorithms related to gaze estimation on HMDs, beginning with eye region segmentation. Semantic segmentation of the eye region is crucial for gaze estimation and is often the starting point for many gaze estimation algorithms. Early eye region segmentation algorithms like [84] are mainly driven by iris texture and sclera extraction [51]. Nowadays, there is an increasing shift towards multi-region eye segmentation, which is typically realized through deep learning. Unlike other semantic segmentation tasks, eye region segmentation is often challenging due to low-resolution images and blurs.
The encoder-decoder network called EyeNet [48] addresses these issues by using residual blocks in both encoder and decoder to improve gradient flow and using convolutional block attention modules (CBAM [85]) to enhance boundary sharpness and accuracy, accomplishing a total score of 0.97 on the EDS evaluation metric. [49] explored the applicability of CycleGANs [86] for eye segmentation and suggested three different generative adversarial networks (GANs) for segmentation, image refinement, and image generation, respectively, that were trained with cyclic loss to prevent the discriminator from overfitting. Conventionally, eye segmentation is vulnerable to occlusion caused by eyelids and eyelashes. [50] addressed this challenge with a segmentation framework, EllSeg, that is robust against occlusions, which can be implemented jointly with other pupil and iris ellipse segmentation methods in an encoder-decoder pattern.
@@
Computational resources, such as computational power, memory, and storage, are often bottlenecks of HMDs. [52] alleviated the computational burden on embedded systems by introducing their multi-scaled segmentation network (Eye-MMS) based on a multi-scale interconnected CNN. This approach achieved a 3% loss in accuracy compared to the original model while successfully reducing the number of model parameters from 6,574k to 80k. [53] proposed a lightweight encoder-decoder segmentation network, EyeSeg, which adopts a customized loss function to tackle labeled data shortage. EyeSeg contains only 190k parameters while yielding a 94.5% mean intersection over union (mIOU) on the OpenEDS dataset. Another lightweight segmentation network, RITnet [54], combines U-Net [87] and DenseNet [88], which is only 0.98 MB and allows eye segmentation at 300 Hz in real-time, with an mIOU of 95.3%, on the OpenEDS dataset.
@@
Another challenge in eye region segmentation is the lack of high-quality eye images. [55] overcame this by generating eye images from semantic segmentation with their novel model called D-ID-NET. The pipeline of this model consists of two phases: in the first phase, a domain network (D-Net) synthesizes identity-irrelevant images from semantic labels, and in the second phase, identity-specific information is introduced into the images by an identity-specific network (ID-Net). Both networks are CNNs with the same structure but different training strategies. [56] implemented a similar concept with their model Seg2Eye, which generates content-preserving eye images from semantic segmentation, resembling style transfer [89] in the sense that semantic segmentation defines the content of generated images. In contrast, their styles are controlled by style features extracted from images of the target person.
[51] extensively studied domain adaptation for eye segmentation when only a few source images are annotated, while most data are not labeled. They systematically investigated the impact of annotated data by training the model in supervised, unsupervised, and semi-supervised manners, respectively, and varying the amount of labeled eye images in the target domain during training. More recently, an image-to-image translation network, GeoMaskGAN [57], was introduced, and it accounts for geometric consistency and consumes a pair of eye images and an eye segmentation mask as input, and outputs a new pair while reducing the geometric gap between translated images and original ones.